During a rainstorm, the rain is flowing into a large holding tank. At the rate the rain is falling, the empty tank would fill in 8 hours. At the bottom of the tank is a spigot to despense water. Typically, it takes about 12 hours with the spigot wide open to empty the water in a full tank. If the tank is empty and the spigot has been accidentally left open, and the rain falls at the constant rate, how long would it take for the tank to fill completely?
Answer by stanbon(75887)

During a rainstorm, the rain is flowing into a large holding tank. At the rate the rain is falling, the empty tank would fill in 8 hours.
At the bottom of the tank is a spigot to despense water.
Typically, it takes about 12 hours with the spigot wide open to empty the water in a full tank.
If the tank is empty and the spigot has been accidentally left open, and the rain falls at the constant rate, how long would it take for the tank to fill completely?
---
Filling rate: 1/8 job/hr
Emptying rate: 1/12 job/hr
Together rate: 1/x job/hr
------------------------------
Equation:
1/8 - 1/12 = 1/x
3x - 2x = 12
x = 12 hrs (time to fill the tank)
=======================================
Cheers,
Stan H.